When Mamma’s Pies decreases the price of pies from R40 to R35, the quantity of the pies 

demanded increases from 800 to 1 150. Calculate the price elasticity of demand for pies at 

Mamma’s Pies for this price range. 

[1] 2,43 

[2] 2,69 

[3] 2,71 

[4] 3,00 

37. When unemployment in South Africa increases, motor vehicles sales usually declines. This fact 

indicates that the __________elasticity of demand for motor vehicles is _________. 

[1] income; positive 

[2] income; negative 

[3] cross; positive 

[4] cross; negative 

38. A seller in a perfectly competitive market faces a demand curve that is 

[1] perfectly inelastic. 

[2] perfectly elastic. 

[3] relatively inelastic. 

[4] relatively elastic.


1
Expert's answer
2021-02-04T14:43:14-0500

The price elasticity of demand for pies at Mamma’s Pies for this price range is:

"Ed = \frac{1150 - 800} {35 - 40} ×\frac{35 + 40} {1150 + 800} = 2.69."

So, the correct answer is [2] 2,69.

37. This fact indicates that the income elasticity of demand for motor vehicles is positive.

So, the correct answer is [1] income; positive.

38. A seller in a perfectly competitive market faces a demand curve that is perfectly elastic.

So, the correct answer is [2] perfectly elastic.